# Astrophysical Consequences of Black Holes in Asymptotically Safe Quantum Gravity: Shadows and Strong Gravitational Lensing

### v2

@@ IV.4 Time delay in the strong field limit
Time delay is one of the most significant observables in the strong gravitational lensing effect. It arises from the time difference between the formation of two relativistic images. This time difference occurs because the photon traverses different paths around the black hole. Since the travel time for photon paths associated with distinct relativistic images is different, a time difference exists between these images. If the time signals from two relativistic images are obtained through observations, one can calculate the time delay between the two signals [54, 128]. The time taken by a photon to orbit the black hole is given by [128]
@@
Using the above equation, the time delays for a Schwarzschild black hole ($S_{0}=S_{2}=0$) and the asymptotically safe quantum gravity black hole, with $S_{0}=S_{2}=0.1$ and $m_{0}=m_{2}=1$ have been estimated in Table 4 for several supermassive black holes located in the centers of nearby galaxies. The deviation of the time delays between the first and second relativistic images, for the asymptotically safe quantum gravity black hole from Schwarzschild black hole $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{ASQG}$minutes with $S_{0}=S_{2}=0.1$ and $m_{0}=m_{2}=1$ has been displayed for the cases of the supermassive black holes $M87^{*}$ and $SgrA^{*}$ in Fig. 12. Fig. 12 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, $\delta\Delta T_{2,1}$ increases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, decreases the value of the observable quantity $\delta\Delta T_{2,1}$.
Figure 12: The deviation of the time delays between the first and second relativistic images, for the asymptotically safe quantum gravity black hole from Schwarzschild black hole $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{ASGQ}$minutes is presented as a function of the parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$.
Galaxy $M(M_{\odot})$ $D_{ol}(Mpc)$ $\Delta T_{2,1}$ SchwarzschildBH ($S_{0}=S_{2}=0$) $\Delta T_{2,1}$ ASQG BH($S_{0}=S_{2}=0.1$,$m_{0}=m_{2}=1$)
@@
Table 4: Estimation of the time delay for various supermassive BHs for the Schwarzschild metric and for the asymptotically safe quantum gravity black holes, respectively. The masses and the distances are expressed in Solar Masses and Mpc units [129]. The time delays $\Delta T_{2,1}$ are estimated in minutes.
@@ IV.5 Results and discussions
Using observations of strong gravitational fields, we have applied the methodology introduced by Bozza [56] to distinguish between various types of spherically symmetric black holes and investigate their astrophysical implications, particularly for supermassive black holes located at the centers of nearby galaxies. This methodology has been extensively employed to test the key properties of asymptotically safe quantum gravity black holes and compare them with the standard Schwarzschild general relativity solution. In our analysis, we investigated several key aspects of astrophysical relevance. First, we computed the strong lensing coefficients $\bar{a}$, $\bar{b}$, and $u_{ph}$ (see Table 2). Utilizing these coefficients, we derived the strong deflection angle along with several observable quantities, such as the angular position of the innermost image $\theta_{\infty}$, the image separation $X$, and the magnification ratio $r_{mag}$, specifically for the supermassive black holes M87 and Sgr A (see Table 3). Furthermore, we examined the time delay between relativistic images (see Table 4) for multiple supermassive black holes located at the centers of nearby galaxies.
Angular position $\theta_{\infty}$ for the set of images.
We have numerically determined the angular positions of the relativistic images for various supermassive black holes within the asymptotically safe quantum gravity framework. The results for the supermassive black holes M87 and Sgr A are summarized in Table 5.
@@
Table 5: Admissible range for the angular position of the innermost relativistic images $\theta_{\infty}$, is analyzed for the supermassive black holes M87 and Sgr A. This range is determined for varying values of the asymptotically safe quantum gravity black hole parameters $S_{2}$, $m_{0}$, and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As one can see from Tables 2 and 5, the angular position of the innermost relativistic images $\theta_{\infty}$ depends on the values of the asymptotically safe quantum gravity black hole parameters $S_{2}$, $m_{0}$, and $m_{2}$. An increase in the value of $m_{2}$ from 0.3 to 1.1 will lead, for the black holes $M87^{*}$ and $SgrA^{*}$, to an increase in $\theta_{\infty}$ for fixed values of the remaining parameters, while $\theta_{\infty}$ decreases with an increase in the parameter $S_{2}$ from 0.1 to 0.3, assuming the other parameters are fixed. Furthermore, considering the same mass and distance of the black hole, it is observed that the angular position of the innermost relativistic images $\theta_{\infty}$ for the asymptotically safe quantum gravity black hole is smaller than that in the case of a Schwarzschild black hole. Consequently, the fundamental asymptotically safe theory of quantum gravity has a significant effect on the potentially observable angular positions of the images $\theta_{\infty}$, enabling a distinction between these images and those associated with other classical black holes.
The angular separation.
We have numerically obtained the angular separation between the outermost and innermost relativistic images $X$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ within the framework of asymptotically safe quantum gravity, as summarized in Table 6.
@@
Table 6: Admissible range for the angular separation between outermost and innermost relativistic images $X$, is analyzed for the supermassive black holes M87 and Sgr A. This range is determined for varying values of the asymptotically safe quantum gravity parameters $S_{2}$, $m_{0}$, and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As one can see from Tables 2 and 6, the angular separation between outermost and innermost relativistic images $X$ depends on the values of the asymptotically safe quantum gravity black hole parameters $S_{2}$, $m_{0}$, and $m_{2}$. An increase in the value of $m_{2}$ from 0.3 to 1.1 will lead, for the black holes $M87^{*}$ and $SgrA^{*}$, to a decrease in $X$ for fixed values of the remaining parameters, while $X$ increases with an increase in the parameter $S_{2}$ from 0.1 to 0.3, assuming the other parameters are fixed. Furthermore, considering the same mass and distance of the black hole, it is observed that the angular separation between outermost and innermost relativistic images $X$ for the asymptotically safe quantum gravity black hole is larger than that in the case of a Schwarzschild black hole. Consequently, the fundamental asymptotically safe theory of quantum gravity has a significant effect on the potentially observable angular separation between the outermost and innermost relativistic images $X$, enabling a distinction between these images and those associated with other classical black holes.
Relative magnification.
We have numerically obtained the relative magnification of the relativistic images $r_{mag}$for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ within the framework of asymptotically safe quantum gravity, as summarized in Table 7.
@@
Table 7: Admissible range for the relative magnification of the relativistic images $r_{mag}$, is analyzed for the supermassive black holes M87 and Sgr A. This range is determined for varying values of the asymptotically safe quantum gravity parameters $S_{2}$, $m_{0}$, and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As illustrated in Tables 2 and 7, the relative magnification of the relativistic images, denoted as $r_{mag}$, is influenced by the parameters of the asymptotically safe quantum gravity black hole, specifically $S_{2}$, $m_{0}$, and $m_{2}$. Increasing the value of $m_{2}$ from 0.3 to 1.1 results in a corresponding rise in $r_{mag}$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$, assuming the other parameters remain constant. Conversely, $r_{mag}$ decreases as the parameter $S_{2}$ increases from 0.1 to 0.3, again with other parameters held fixed.
Moreover, when comparing black holes of the same mass and distance, it is found that the relative magnification $r_{mag}$ for the asymptotically safe quantum gravity black hole is less than that for a Schwarzschild black hole. This indicates that the fundamental aspects of asymptotically safe quantum gravity significantly affect the observable relative magnification of the relativistic images, allowing for differentiation between these images and those linked to other classical black holes. It is also important to note that the relative magnification of the relativistic images is independent of the black hole’s mass or distance.
The Einstein rings.
@@
We have also evaluated the time delays $\Delta T_{2,1}$ between the first and second-order relativistic images for various supermassive black holes within the framework of the asymptotically safe quantum gravity black hole, using the parameter values $(S_{0}=S_{2}=0.1,\,m_{0}=m_{2}=1)$, as well as for the Schwarzschild black hole $(S_{0}=S_{2}=0)$. The time delay $\Delta T_{2,1}$ for the asymptotically safe quantum gravity black hole (e.g., $\sim 23115.7$ minutes for NGC $3842$) is significantly smaller compared to the Schwarzschild black hole (e.g., $\sim 24303.7$ minutes for NGC $3842$). Thus, the deviation in the time delays for the asymptotically safe quantum gravity black hole with $(S_{0}=S_{2}=0.1,\,m_{0}=m_{2}=1)$ from the standard Schwarzschild black hole $(S_{0}=S_{2}=0)$ is $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{ASQG}=1188\,\text{minutes}.$ This suggests that if the first and second relativistic images are distinguishable, the time delay between them could serve as a valuable indicator to differentiate the asymptotically safe quantum gravity black hole from the classical Schwarzschild black hole.
V Conclusions, and final remarks
@@
In the present paper, we have discussed the shadow and strong gravitational lensing phenomena by the asymptotically safe quantum gravity black holes, described by a solution obtained by the quantum effective action and the respective quantum equations of motion from multi-graviton correlation functions in asymptotically safe quantum gravity [101]. A black hole in asymptotically safe quantum gravity is described by five parameters, with $M$ representing the mass of the black hole. The two free parameters $S_{0}$ and $S_{2}$ characterize the strength of the exponentially decaying Yukawa corrections. Additionally, the masses $m_{0}$ and $m_{2}$ of the spin-0 and spin-2 modes are associated with various curvature terms’ couplings. This has been discussed in detail in [101]. The classical Schwarzschild black hole is obtained when $S_{0}=S_{2}=0$. The present study highlights that traditional black hole solutions derived from classical general relativity may not fully capture the complexities of black holes in the context of quantum gravity. The asymptotic safety scenario suggests that gravitational interactions approach a fixed point at high energy scales, indicating new physics beyond classical theories. Quantum black hole solutions exhibit significant deviations from classical predictions, particularly outside the event horizon. These deviations could manifest in gravitational lensing phenomena, where the curvature of spacetime influences the paths of light rays passing near massive objects. In this study, we primarily restrict our analysis to positive values of the black hole parameters $S_{0}$,$S_{2}$,$m_{0}$, and $m_{2}$. The study of light deflection (gravitational lensing) in the strong field limit, particularly in regions dominated by the gravitational effects of asymptotically safe quantum gravity black holes, offers powerful theoretical and observational tools for distinguishing between standard general relativistic models and modified gravity models. Although the radial coordinate dependence of the asymptotically safe quantum gravity metric takes a fixed form, when interpreted as a galactic metric, the bending angle of light and other optical effects are significantly influenced by the metric parameters, as well as by the baryonic mass and radius of the supermassive black hole. The asymptotically safe quantum gravity metric dictates the specific values of the bending angles and other lensing effects, and these values differ substantially from those predicted by the classical Schwarzschild metric. For example, when $S_{0}=0.1$,$S_{2}=0.1$,$m_{0}=1$, and $m_{2}=1$, the time delay of the photons is much smaller in the asymptotically safe quantum gravity than in standard general relativity. This may be related to the gravitational light deflection angle being larger than the value predicted by the standard general relativistic approach. We observe significant differences in the lensing effects when comparing our results for the asymptotically safe quantum gravity black holes with those of standard Schwarzschild black holes. As a result, the asymptotically safe quantum gravity black hole, influenced by moving fluids and sound waves, can be detected more easily and distinguished from ordinary astrophysical black holes, such as the Schwarzschild black hole. Consequently, the study of gravitational lensing may provide evidence for the existence of asymptotically safe quantum gravity effects in the Universe. Our findings represent a promising and significant step toward connecting the fundamental theory of asymptotically safe quantum gravity to potentially feasible black hole spacetimes.

### v3

@@ IV.4 Time delay in the strong field limit
Time delay is one of the most significant observables in the strong gravitational lensing effect. This arises from the time difference between the formation of two relativistic images. This time difference occurs because the photon traverses different paths around the black hole. Since the travel time for photon paths associated with distinct relativistic images is different, a time difference exists between these images. If the time signals from two relativistic images are obtained through observations, one can calculate the time delay between the two signals [63, 129]. The time taken by a photon to orbit the black hole is given by [129]
@@
Using the above equation, the time delays for a Schwarzschild black hole ($S_{0}=S_{2}=0$) and the quadratic gravity black hole, with $S_{0}=S_{2}=0.1$ and $m_{0}=m_{2}=1$ have been estimated in Table 6 for several supermassive black holes located in the centers of nearby galaxies. The deviation of the time delays between the first and second relativistic images, for the quadratic gravity black hole from the Schwarzschild black hole $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{ASquadraticgravity}$minutes with $S_{0}=S_{2}=0.1$ and $m_{0}=m_{2}=1$ has been displayed for the cases of supermassive black holes $M87^{*}$ and $SgrA^{*}$ in Fig. 14. Fig. 14 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, $\delta\Delta T_{2,1}$ increases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, decreases the value of the observable quantity $\delta\Delta T_{2,1}$.
Figure 14: The deviation of the time delays between the first and second relativistic images, for the quadratic gravity black hole from Schwarzschild black hole $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{QG}$ minutes is presented as a function of the parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$.
Galaxy $M(M_{\odot})$ $D_{ol}(Mpc)$ $\Delta T_{2,1}$ SchwarzschildBH ($S_{0}=S_{2}=0$) $\Delta T_{2,1}$ quadratic gravity BH($S_{0}=S_{2}=0.1$,$m_{0}=m_{2}=1$)
@@
Table 6: Estimation of the time delay for various supermassive BHs for the black hole-like spacetime in quadratic gravity and the Schwarzschild spacetime, respectively. The masses and distances are given in solar masses and megaparsecs (Mpc), respectively [130]. The time delays $\Delta T_{2,1}$ are estimated in minutes.
@@ IV.5 Results and discussions
Using observations of strong gravitational fields, we employed the methodology introduced by Bozza [65] to differentiate between various types of spherically symmetric black holes and explore their astrophysical consequences, especially for supermassive black holes at the centers of nearby galaxies. This approach has been widely used to test the fundamental properties of black holes in quadratic gravity, comparing them with the standard Schwarzschild solution in general relativity. In our analysis, we examined several key aspects of astrophysical significance, including the calculation of strong lensing coefficients $\bar{a}$, $\bar{b}$ and $u_{ph}$ (see Table 4). Using these coefficients, we derived the strong deflection angle along with several observable quantities, such as the angular position of the innermost image $\theta_{\infty}$, the image separation $X$ and the magnification ratio $r_{mag}$, specifically for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ (see Table 5). Furthermore, we examined the time delay between relativistic images (see Table 6) for multiple supermassive black holes located at the centers of nearby galaxies.
Angular position $\theta_{\infty}$ for the set of images:
We have numerically determined the angular positions of the relativistic images for various supermassive black holes within the quadratic gravity framework. The results for supermassive black holes $M87^{*}$ and $SgrA^{*}$ are summarized in Table 7.
@@
Table 7: Admissible range for the angular position of the innermost relativistic images $\theta_{\infty}$, is analyzed for the supermassive black holes $M87^{*}$ and $SgrA^{*}$. This range is determined for varying values of the quadratic gravity black hole parameters $S_{2}$, $m_{0}$ and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As one can see from Tables 4 and 7, the angular position of the innermost relativistic images $\theta_{\infty}$ depends on the values of the quadratic gravity black hole parameters $S_{2}$, $m_{0}$ and $m_{2}$. An increase in the value of $m_{2}$ from 0.3 to 1.1 will lead, for the black holes $M87^{*}$ and $SgrA^{*}$, to an increase in $\theta_{\infty}$ for fixed values of the remaining parameters, while $\theta_{\infty}$ decreases with an increase in the parameter $S_{2}$ from 0.1 to 0.3, assuming the other parameters are fixed. Furthermore, considering the same mass and distance of the black hole, it is observed that the angular position of the innermost relativistic images $\theta_{\infty}$ for the quadratic gravity black hole is smaller than that in the case of a Schwarzschild black hole. Consequently, fundamental quadratic gravity has a significant effect on the potentially observable angular positions of the images $\theta_{\infty}$, allowing a distinction between these images and those associated with other classical black holes.
The angular separation:
We have numerically obtained the angular separation between the outermost and innermost relativistic images $X$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ within the framework of quadratic gravity, as summarized in Table 8.
@@
Table 8: Admissible range for the angular separation between outermost and innermost relativistic images $X$, is analyzed for the supermassive black holes $M87^{*}$ and $SgrA^{*}$. This range is determined for varying values of the quadratic gravity parameters $S_{2}$, $m_{0}$ and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As one can see from Tables 4 and 8, the angular separation between the outermost and innermost relativistic images $X$ depends on the values of the quadratic gravity black hole parameters $S_{2}$, $m_{0}$ and $m_{2}$. An increase in the value of $m_{2}$ from 0.3 to 1.1 will lead, for the black holes $M87^{*}$ and $SgrA^{*}$, to a decrease in $X$ for fixed values of the remaining parameters, while $X$ increases with an increase in the parameter $S_{2}$ from 0.1 to 0.3, assuming the other parameters are fixed. Furthermore, considering the same mass and distance of the black hole, it is observed that the angular separation between outermost and innermost relativistic images $X$ for the quadratic gravity black hole is larger than that in the case of a Schwarzschild black hole. Consequently, fundamental quadratic gravity has a significant effect on the potentially observable angular separation between the outermost and innermost relativistic images $X$, allowing a distinction between these images and those associated with other classical black holes.
Relative magnification:
We have numerically obtained the relative magnification of the relativistic images $r_{mag}$for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ within the framework of quadratic gravity, as summarized in Table 9.
@@
Table 9: Admissible range for the relative magnification of the relativistic images $r_{mag}$, is analyzed for the supermassive black holes $M87^{*}$ and $SgrA^{*}$. This range is determined for varying values of the quadratic gravity parameters $S_{2}$, $m_{0}$ and $m_{2}$ while keeping the parameter $S_{0}$ constant.
As illustrated in Tables 4 and 9, the relative magnification of the relativistic images, denoted as $r_{mag}$, is influenced by the parameters of the quadratic gravity black hole, specifically $S_{2}$, $m_{0}$ and $m_{2}$. Increasing the value of $m_{2}$ from 0.3 to 1.1 results in a corresponding rise in $r_{mag}$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$, assuming the other parameters remain constant. Conversely, $r_{mag}$ decreases as the parameter $S_{2}$ increases from 0.1 to 0.3, again with other parameters held fixed.
Moreover, when comparing black holes of the same mass and distance, it is found that the relative magnification $r_{mag}$ for the quadratic gravity black hole is less than that for a Schwarzschild black hole. This indicates that the fundamental aspects of quadratic gravity significantly affect the observable relative magnification of the relativistic images, allowing for differentiation between these images and those linked to other classical black holes. It is also important to note that the relative magnification of the relativistic images is independent of the black hole’s mass or distance.
The Einstein rings:
@@
We have also evaluated the time delays $\Delta T_{2,1}$ between the first and second-order relativistic images for various supermassive black holes within the framework of the quadratic gravity black hole, using the parameter values $(S_{0}=S_{2}=0.1,\,m_{0}=m_{2}=1)$, as well as for the Schwarzschild black hole $(S_{0}=S_{2}=0)$. The time delay $\Delta T_{2,1}$ for the quadratic gravity black hole (e.g., $\sim 23115.7$ minutes for NGC $3842$) is significantly smaller compared to the Schwarzschild black hole (e.g., $\sim 24303.7$ minutes for NGC $3842$). Thus, the deviation in the time delays for the quadratic gravity black hole with $(S_{0}=S_{2}=0.1,\,m_{0}=m_{2}=1)$ from the standard Schwarzschild black hole $(S_{0}=S_{2}=0)$ is $\delta\Delta T_{2,1}=\Delta T_{2,1}^{Sch}-\Delta T_{2,1}^{QG}=1188,\text{minutes}.$ This suggests that if the first and second relativistic images are distinguishable, the time delay between them could serve as a valuable indicator to differentiate the quadratic gravity black hole from the classical Schwarzschild black hole.
V Conclusions and final remarks
This paper explores the shadow and strong gravitational lensing phenomena associated with black holes in quadratic gravity. These black holes are described by a solution derived from the quantum effective action and the corresponding quantum equations of motion, which are obtained from multi-graviton correlation functions in quadratic gravity [102]. A black hole in quadratic gravity is described by five parameters, with $M$ representing the mass of the black hole. The two free parameters $S_{0}$ and $S_{2}$ characterize the strength of the exponentially decaying Yukawa corrections. Additionally, the masses $m_{0}$ and $m_{2}$ of the spin-0 and spin-2 modes are associated with various couplings of curvature terms. This has been discussed in detail in [102].
Equations (2)-(4) provide the parametrization of asymptotically flat solutions within the linearized regime. These equations ensure that the asymptotically flat solutions extend smoothly to $r\to\infty$. By systematically varying the parameters, one can identify a variety of topologically distinct solutions, each characterized by specific scaling behaviors near $r=0$ or, in some cases, by a termination point $r_{\text{term}}>0$. The defining features of these solution classes (Type Ia, Type Ib, Type Ic, Type II and Type III) emerge from the distinct initial conditions outlined in Table 2 of the referenced literature [24]. From the perspective of quadratic gravity, the geometry described by Eqs. (2)-(4) represents the gravitational field in the linear approximation. However, this framework breaks down in the strong gravity regime. In this regime, the event horizon becomes sensitive to exponentially small Yukawa corrections and non-zero values of $S_{0}$ and $S_{2}$ can give rise to wormholes and naked singularities. Consequently, the standard formulas used for shadow imaging must be applied with caution, as new photon trajectories can emerge that would otherwise be blocked by an event horizon. A detailed analysis of this phenomenon is available in the literature [24]. Quadratic gravity introduces significant effects on Event Horizon Telescope (EHT) observations, with profound implications across multiple domains. The Schwarzschild-like geometry described by Eqs. (2)-(4) is discovered to be part of a diverse solution space that includes horizonless configurations, naked singularities and wormhole structures. The inclusion of Yukawa corrections and additional parameters like $S_{0}$ and $S_{2}$ opens up new photon trajectories that do not occur in standard black hole models. Consequently, traditional shadow imaging methods must be carefully reconsidered as these new pathways alter the observed shadow properties. A significant observation is the impact of such modifications on the photon ring and shadow size, particularly for TypeIc and TypeII geometries, which exhibit distinguishable features in their intensity profiles (see [24] for more details). These findings emphasize that shadow observations using the EHT can effectively constrain the parameter space of quadratic gravity. Their findings lead to two significant implications. First, the EHT demonstrates the ability to explore and even exclude certain regions of the phase space associated with static, spherically symmetric and asymptotically flat vacuum solutions in quadratic gravity. Second, this phase space includes horizonless naked singularities and wormhole geometries with intensity profiles nearly identical to those produced by the Schwarzschild geometry. Consequently, based on these observations alone, it remains challenging to determine whether the imaged object genuinely has an event horizon. In the present paper, we have constrained the black hole parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$ through shadow observables (the angular diameter of the black hole shadow) using the EHT observation data for the supermassive black holes M87* and Sgr A*. This was done for both cases: the new photon sphere radius $r^{L}_{ph}$ and the standard photon sphere radius $r_{ph}$. It is observed that, in both cases, the obtained results fall within a finite parameter space and are consistent with the EHT observation data for M87* and Sgr A*. Our work interprets the EHT observational results for black-hole-type solutions within the framework of a clear quantum gravity perspective, grounded in observational data. Thus, the Schwarzschild black hole solution in quadratic gravity emerges as a viable model.
@@ V Conclusions and final remarks
The classical Schwarzschild black hole is obtained when $S_{0}=S_{2}=0$. The present study highlights that traditional black hole solutions derived from classical general relativity may not fully capture the complexities of black holes in the context of quantum gravity. The asymptotic safety scenario suggests that gravitational interactions approach a fixed point at high-energy scales, indicating new physics beyond classical theories. Quantum black hole solutions in quadratic gravity exhibit significant deviations from classical predictions, particularly outside the event horizon. These deviations could manifest in gravitational lensing phenomena, where the curvature of spacetime influences the paths of light rays passing near massive objects. In this study, we primarily restrict our analysis to positive values of the black hole parameters $S_{0}$,$S_{2}$,$m_{0}$ and $m_{2}$. The study of light deflection (gravitational lensing) in the strong field limit, particularly in regions dominated by the gravitational effects of quadratic gravity black holes, offers powerful theoretical and observational tools for distinguishing between standard general relativistic models and modified gravity models. Although the radial coordinate dependence of the quadratic gravity metric takes a fixed form, when interpreted as a galactic metric, the bending angle of light and other optical effects are significantly influenced by the metric parameters, as well as by the baryonic mass and radius of the supermassive black hole. The quadratic gravity metric dictates the specific values of the bending angles and other lensing effects, and these values differ substantially from those predicted by the classical Schwarzschild metric.
For example, when $S_{0}=0.1$,$S_{2}=0.1$,$m_{0}=1$, and $m_{2}=1$, the time delay of the photons is much smaller in quadratic gravity than in standard general relativity. This may be related to the gravitational light deflection angle being larger than the value predicted by the standard general relativistic approach. We observe significant differences in the lensing effects when comparing our results for the quadratic gravity black holes with those of standard Schwarzschild black holes.
Moreover, the study of black hole-like solutions in quadratic gravity can be distinguished from standard astrophysical black holes, such as the Schwarzschild black holes($S_{0}=0$,$S_{2}=0$), through shadow and strong gravitational lensing observations.
As a result, the quadratic gravity black hole, influenced by moving fluids and sound waves, can be detected more easily and distinguished from ordinary astrophysical black holes, such as the Schwarzschild black hole. Consequently, the study of gravitational lensing may provide evidence for the existence of quadratic gravity effects in the Universe. Our findings represent a promising and significant step toward connecting the fundamental theory of quadratic gravity to potentially feasible black hole spacetimes.